A person drops a cylindrical steel bar ( Y = 1.700 × 10 11 Pa ) from a height of 3.70 m (distance between the floor and the bottom of the vertically oriented bar). The bar, of length L = 0.950 m , radius R = 0.00650 m , and mass m = 2.000 kg , hits the floor and bounces up, maintaining its vertical orientation. Assuming the collision with the floor is elastic, and that no rotation occurs, what is the maximum compression of the bar? maximum compression: m
Answer #1

Given that, Youngs modulus of steel bar is , Y = 1-7X10Pa Height, h= 3.70m length, L=0.950m radius, R=0.00 650m mass, m=2.0As in the case of elastic collision energy remains conserved. Ex=Ep. Substitute the values, mas = Į (T (HRT)) yn or, y = /2 m